Find the distance between the two points (3,-4) and (-4,5)
Answer by mukhopadhyay(490)

Use the formula d^2=(x2-x1)^2 + (y2-y1)^2
So, d^2 = [3-(-4)]^2 + [(-4)-5]^2
=>d^2 = 7^2 + 9^2
=>d^2 = 49+81
=>d^2 = 130
=>d = sqrt(130)
The distance between two given points is sqrt(130) units